Sally wants to know how much water to add to a 40 ml of 70% acid to make a concentration of 50% acid.
Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12844)

.70*40=.50(40+x)
28=20+.5x
.5x=28-20
.5x=8
x=8/.5
x=16 ml of water is needed.
Proof:
.70*40=.50(40+16)
28=.5*56
28=28

Answer by stanbon(75887)

Sally wants to know how much water to add to a 40 ml of 70% acid to make a concentration of 50% acid.
---------------------------------
Equation:
acid + acid = acid
0.70*40 + 0*x = 0.50(x+40)
Multiply thru by 100 to get:
70*40 + 0 = 40x + 50*40
40x = 20*40
x = 20 ml (amount of water to add)
=======================================
Cheers,
Stan H.